# Global End-Effector Pose Control of an Underactuated Aerial Manipulator via Reinforcement Learning

### v1

@@ Abstract
Aerial manipulators, which combine robotic arms with multi-rotor drones, face strict constraints on arm weight and mechanical complexity. In this work, we study a lightweight 2-degree-of-freedom (DoF) arm mounted on a quadrotor via a differential mechanism, capable of full six-DoF end-effector pose control. While the minimal design enables simplicity and reduced payload, it also introduces challenges such as underactuation and sensitivity to external disturbances. To address these, we employ reinforcement learning, training a Proximal Policy Optimization (PPO) agent in simulation to generate feedforward commands for quadrotor acceleration and body rates, along with joint angle targets. These commands are tracked by an incremental nonlinear dynamic inversion (INDI) attitude controller and a PID joint controller, respectively. Flight experiments demonstrate centimeter-level position accuracy and degree-level orientation precision, with robust performance under external force disturbances—including manipulation of heavy loads and pushing tasks. The results highlight the potential of learning-based control strategies for enabling contact-rich aerial manipulation using simple, lightweight platforms.
@@ 1 Introduction
Many aerial manipulators employ redundant arms with multiple degrees of freedom, or add additional tilted / tiltable rotors to achieve full actuation of the base, but at the cost of additional weight and structural complexity (e.g.,  [kim2013aerial, tognon2019truly, afifi2022toward]). Unlike ground and underwater robots, aerial robots must prioritize lightweight design due to their limited thrust-to-weight ratio and restricted onboard energy resources. To address this, we adopt a minimal yet effective design: the Differential Shoulder Aerial Manipulator (DSAM), where a single rigid arm with an end-effector (e.g., a gripper) is connected to the quadrotor base through a differential driven by two rotational actuators, enabling pitch and roll motions (Figure 1). Prior work has shown that a configuration for AM like DSAM is sufficient to achieve global 6-DoF end-effector pose regulation, despite its mechanical simplicity [welde2021dynamically]. However, the problem of designing robust whole-body control algorithms for DSAM remains largely unexplored.
Before developing model-based control methods, we first investigate reinforcement learning (RL) for the whole-body control of DSAM. RL-based controllers offer significant advantages by alleviating the need for analytical models and their simplifications, implicitly considering the system actuation constraints through explorations, enabling fast inference, and supporting adaptation to unmodeled disturbances through domain randomization, all of which are essential for real-world aerial manipulation. To address the sim-to-real gap, we also employ domain randomization during training, and adopt the scheme that hybridizes the RL with downstream low-level controllers [hwangbo2019learning]. Specifically, we employ an incremental nonlinear dynamic inversion (INDI) attitude controller for the quadrotor base, and a PID controller for the arm. We have experimentally validated the proposed algorithm in controlling the 6-DoF end-effector pose, as well as experiments testing the robustness against external forces on the end-effector, including carrying an additional load, and pushing an object.
@@ 2.1 System Design
Aerial manipulators can be grouped into two design families, each aiming to realize end-effector (EE) regulation in SE(3) or a relevant subspace. (i) The first augments an underactuated unidirectional multirotor base with a serial or parallel arm (e.g., [kim2013aerial, chen2025aerial]). (ii) The second employs overactuated drone bases with tilted or tiltable rotors, allowing EEs fixed to the base to realize arbitrary pose tracking in SE(3) (e.g., [rashad2019port, li2024servo, kamel2018voliro, hui2024passive, tognon2018omnidirectional]). Some designs employ both choices (e.g.,  [tognon2019truly, afifi2022toward, malczyk2023multi]). A more detailed testimony of aerial manipulator designs, and their comparisons, is given in the survey [9462539]. We focus on the first family and adopt the DSAM: a quadrotor base connected to an arm through a 2-DoF differential. DSAM is one of the most minimal designs achieving global 6-DoF end-effector pose control, where the dual rotational actuators enlarge arm torques, thereby boosting load and manipulation capability without increasing weight or complexity. Ref. [welde2021dynamically] concludes that task flatness holds for a DSAM in a special condition, namely, the differential is located at the CoM of the quadrotor base. In practice, however, this condition may not hold due to design constraints and CoM variations. In addition, directly designing the control law through differential flatness cannot effectively satisfy input and state constraints, which are essential for aerial robotic manipulators.
@@ 2.2 Planning and Control
Specifically for DSAM, a controller still needs to be designed to reliably track setpoints and trajectories across the full SE(3) task space. Most research on aerial manipulator control employ model-based methods. Reactive control methods, such as feedback linearization and geometric control [lippiello2012exploiting, Zhang_2020, 9345436, 6386021, 6580711], offer computational efficiency but provide limited mechanisms for handling actuator bounds and state constraints. Several works employ model predictive control (MPC) and Model Predictive Path Integral (MPPI) control for aerial manipulators, which improve constraint handling and can reason over short horizons, yet they incur substantial online computation and require careful contact and cost-shaping design in contact-rich tasks [7139850, marti2021full, 9830871]. Motivated by promising progress in robot learning for manipulation (e.g., [rajeswaran2017learning]), recent studies have begun to explore reinforcement learning for aerial manipulators, demonstrating promising results on platforms with fully-actuated bases (e.g., omnidirectional bases with fixed EEs for door opening [cuniato2023learning]). Building on this trend, we present the first whole-body controller for DSAM, based on learning-based method, that globally tracks SE(3) setpoints and addresses the underactuation of the DSAM while embracing its hardware design simplicity. This work will pave the way to future research for DSAM in contact-rich aerial manipulation tasks.
@@ 3 Preliminaries
The system evolves according to coupled nonlinear dynamics given by generalized Lagrangian method presented in [From2012SingularityFreePartTwo]
@@ 4.1.1 Reinforcement Learning Policy
It receives observations from the environment and generates actions for subsequent controllers at 150 Hz. Mean and Variance statistics from training are used to normalize the observation vector, followed by inference using ONNX Runtime engine [onnxruntime] onboard for shorter inference times.
@@ 4.1.2 Inner-loop Quadrotor Attitude Controller
The inner-loop attitude controller, running at 300 Hz, converts the RL outputs into rotor-speed commands $\omega_{\mathrm{cmd}}$ for the quadrotor base. It first maps the RL-generated acceleration and heading to a desired attitude via thrust vector decomposition, and then combines a tilt-prioritized controller [brescianini2018tilt] with incremental nonlinear dynamic inversion (INDI)[sun2022comparative] to compute rotor speeds, using the RL-provided body rates as feedforward terms. This inner-loop attitude controller has demonstrated robustness to model mismatches in the quadrotor’s rotational dynamics, which is particularly important for DSAM, where arm motion induces torque disturbances on the base. For further details, we refer readers to [sun2022comparative].
@@ 4.2 Training of the RL Policy
We model the control architecture (Fig. 2) in our training environment and PPO [schulman2017proximal] to train our agent to perform end-effector pose control.
@@ 4.2.1 Training Environment
We use Isaac Lab [mittal2023orbit] for training the RL-policy as it extensively utilizes GPUs for both simulation and learning, enabling high data throughput and faster learning. The policy training is implemented using SKRL reinforcement learning library [serrano2023skrl], which provides an integration with Isaac Lab. We model the DSAM system in the training environment. Apart from modeling the control architecture (Section 4.1) with low-level controllers, we also model filters, first-order dynamics of rotors.
PPO offers control over multiple aspects of the algorithm, making hyperparameter selection challenging. We begin with the hyperparameters used in [rudin2022learningwalkminutesusing] as a starting point.
@@
To achieve a working policy, we train for 2 billion environment steps ($4096\times 500,000$) that take 5.5 hours on a consumer grade GPU (NVIDIA RTX 2080 Ti).
@@ 5.1 Experimental Setup
The policy trained using the proposed method is deployed on our 2-DoF DSAM, with a total weight of 860g. The moving part (arm plus gripper) weighs 75 g, actuated by two Dynamixel-XC330 motors [robotis_xc330m288] with embedded PID controllers following joint commands. Our control architecture is developed upon Agilicious [Foehn22science], a modular software framework designed for agile quadrotor flight. The control stack entirely runs on the DSAM’s onboard CPU (Raspberry Pi 5), including onboard inference using ONNX Runtime [onnxruntime]. This results in an average inference time of 0.18 ms.
Experiments are conducted in a 13 $\times$ 7 m facility equipped with a motion capture system that provides ground-truth pose measurements at 100 Hz, which is then fused with an EKF running onboard at 300 Hz to obtain the state estimates. A safety controller that only controls the quadrotor base motion [sun2022comparative] is used for take-off and handles workspace violations by immediately bringing the system to hover state. After take-off, control is switched from the safety controller to the learned policy that actuates the quadrotor to reach the commanded end-effector pose.
@@ 5.5 Path following
As seen in Table 3, the RMSE for the figure-8 path is noticeably larger than for the straight line, reflecting the higher demands of multi-axis coordination compared to single-axis motion. The tracking accuracy is further limited because the policy is conditioned only on the immediate target pose and does not explicitly account for higher-order derivatives such as velocity, acceleration, or jerk. Consequently, the policy was not trained specifically for accurate path tracking but rather for pose regulation. Future work could address this by augmenting the observation space with future waypoints or target velocities, which have been shown to improve path-tracking performance in related learning-based control frameworks (e.g.,  [garg2025dual]). Despite these limitations, the policy is able to follow simple paths while maintaining orientation, demonstrating whole-body coordination emerging from pose control alone.
@@ 6.1.3 Body rate of the quadrotor base
We noticed that including body rate in the observation space is also critical for training for DSAM, which is particularly different from previous works on quadrotors [song2023reaching], and AMs with a fully-actuated base [cuniato2023learning]. We argue that including quadrotor angular rates in the observation significantly improves policy training, as they allow the policy to indirectly infer the effect of torque disturbances induced by the arm. This additional information helps the policy learn the coupled base–arm dynamics especially for DSAM with an underactuated unidirectional base, which is, on the other hand, less critical for omnidirectional platforms or simple quadrotors.
@@ 6.2 Inner-Loop Controller Ablations
While incorporating an inner-loop is a common practice in RL-based control to achieve reliable sim-to-real transfer, determining the appropriate level of control abstraction is non-trivial. In particular, for the quadrotor base of DSAM, we replace the attitude controller with a simple proportional controller that maps the body-rate command to the desired body angular acceleration, which is then passed to the low-level INDI controller. Under this setup, we evaluate the traditional “collective-thrust plus body rate” (CTBR) output of the RL policy for quadrotor flight, following prior work [song2023reaching]. INDI is retained, as it plays a critical role in handling disturbances, as highlighted in previous studies [sun2022comparative, ferede2024end]. The training curves in Fig. 6 show that our setup, where the attitude controller accepts reference acceleration (or equivalently attitude) from the RL policy, achieves consistently higher rewards in both orientation and position tracking.

### v2

@@ Abstract
Aerial manipulators, which combine robotic arms with multi-rotor drones, face strict constraints on arm weight and mechanical complexity. In this work, we study a lightweight 2-degree-of-freedom (DoF) arm mounted on a quadrotor via a differential mechanism, capable of full six-DoF end-effector pose control. While the minimal design enables simplicity and reduced payload, it also introduces challenges such as underactuation and sensitivity to external disturbances. To address these, we employ reinforcement learning, training a Proximal Policy Optimization (PPO) agent in simulation to generate feedforward commands for quadrotor acceleration and body rates, along with joint angle targets. These commands are tracked by an incremental nonlinear dynamic inversion (INDI) attitude controller and a PID joint controller, respectively. Flight experiments demonstrate centimeter-level position accuracy and degree-level orientation precision, with robust performance under external force disturbances—including manipulation of heavy loads and pushing tasks. The results highlight the potential of learning-based control strategies for enabling contact-rich aerial manipulation using simple, lightweight platforms. Videos of the experiment and the method are summarized in https://youtu.be/bWLTPqKcCOA.
@@ 1 Introduction
Many aerial manipulators employ redundant arms with multiple degrees of freedom, or add additional tilted / tiltable rotors to achieve full actuation of the base, but at the cost of additional weight and structural complexity (e.g.,  [16, 35, 1]). Unlike ground and underwater robots, aerial robots must prioritize lightweight design due to their limited thrust-to-weight ratio and restricted onboard energy resources. To address this, we adopt a minimal yet effective design: the Differential Shoulder Aerial Manipulator (DSAM), where a single rigid arm with an end-effector (e.g., a gripper) is connected to the quadrotor base through a differential driven by two rotational actuators, enabling pitch and roll motions (Figure 1). Prior work has shown that a configuration for AM like DSAM is sufficient to achieve global 6-DoF end-effector pose regulation, despite its mechanical simplicity [37]. However, the problem of designing robust whole-body control algorithms for DSAM remains largely unexplored.
Before developing model-based control methods, we first investigate reinforcement learning (RL) for the whole-body control of DSAM. RL-based controllers offer significant advantages by alleviating the need for analytical models and their simplifications, implicitly considering the system actuation constraints through explorations, enabling fast inference, and supporting adaptation to unmodeled disturbances through domain randomization, all of which are essential for real-world aerial manipulation. To address the sim-to-real gap, we also employ domain randomization during training, and adopt the scheme that hybridizes the RL with downstream low-level controllers [14]. Specifically, we employ an incremental nonlinear dynamic inversion (INDI) attitude controller for the quadrotor base, and a PID controller for the arm. We have experimentally validated the proposed algorithm in controlling the 6-DoF end-effector pose, as well as experiments testing the robustness against external forces on the end-effector, including carrying an additional load, and pushing an object.
@@ 2.1 System Design
Aerial manipulators can be grouped into two design families, each aiming to realize end-effector (EE) regulation in SE(3) or a relevant subspace. (i) The first augments an underactuated unidirectional multirotor base with a serial or parallel arm (e.g., [16, 5]). (ii) The second employs overactuated drone bases with tilted or tiltable rotors, allowing EEs fixed to the base to realize arbitrary pose tracking in SE(3) (e.g., [28, 17, 15, 13, 36]). Some designs employ both choices (e.g.,  [35, 1, 21]). A more detailed testimony of aerial manipulator designs, and their comparisons, is given in the survey [25]. We focus on the first family and adopt the DSAM: a quadrotor base connected to an arm through a 2-DoF differential. DSAM is one of the most minimal designs achieving global 6-DoF end-effector pose control, where the dual rotational actuators enlarge arm torques, thereby boosting load and manipulation capability without increasing weight or complexity. Ref. [37] concludes that task flatness holds for a DSAM in a special condition, namely, the differential is located at the CoM of the quadrotor base. In practice, however, this condition may not hold due to design constraints and CoM variations. In addition, directly designing the control law through differential flatness cannot effectively satisfy input and state constraints, which are essential for aerial robotic manipulators.
@@ 2.2 Planning and Control
Specifically for DSAM, a controller still needs to be designed to reliably track setpoints and trajectories across the full SE(3) task space. Most research on aerial manipulator control employ model-based methods. Reactive control methods, such as feedback linearization and geometric control [19, 38, 20, 18, 26], offer computational efficiency but provide limited mechanisms for handling actuator bounds and state constraints. Several works employ model predictive control (MPC) and Model Predictive Path Integral (MPPI) control for aerial manipulators, which improve constraint handling and can reason over short horizons, yet they incur substantial online computation and require careful contact and cost-shaping design in contact-rich tasks [12, 22, 4]. Motivated by promising progress in robot learning for manipulation (e.g., [27]), recent studies have begun to explore reinforcement learning for aerial manipulators, demonstrating promising results on platforms with fully-actuated bases (e.g., omnidirectional bases with fixed EEs for door opening [6]). Building on this trend, we present the first whole-body controller for DSAM, based on learning-based method, that globally tracks SE(3) setpoints and addresses the underactuation of the DSAM while embracing its hardware design simplicity. This work will pave the way to future research for DSAM in contact-rich aerial manipulation tasks.
@@ 3 Preliminaries
The system evolves according to coupled nonlinear dynamics given by generalized Lagrangian method presented in [10]
@@ 4.1.1 Reinforcement Learning Policy
It receives observations from the environment and generates actions for subsequent controllers at 150 Hz. Mean and Variance statistics from training are used to normalize the observation vector, followed by inference using ONNX Runtime engine [7] onboard for shorter inference times.
@@ 4.1.2 Inner-loop Quadrotor Attitude Controller
The inner-loop attitude controller, running at 300 Hz, converts the RL outputs into rotor-speed commands $\omega_{\mathrm{cmd}}$ for the quadrotor base. It first maps the RL-generated acceleration and heading to a desired attitude via thrust vector decomposition, and then combines a tilt-prioritized controller [3] with incremental nonlinear dynamic inversion (INDI)[34] to compute rotor speeds, using the RL-provided body rates as feedforward terms. This inner-loop attitude controller has demonstrated robustness to model mismatches in the quadrotor’s rotational dynamics, which is particularly important for DSAM, where arm motion induces torque disturbances on the base. For further details, we refer readers to [34].
@@ 4.2 Training of the RL Policy
We model the control architecture (Fig. 2) in our training environment and PPO [31] to train our agent to perform end-effector pose control.
@@ 4.2.1 Training Environment
We use Isaac Lab [24] for training the RL-policy as it extensively utilizes GPUs for both simulation and learning, enabling high data throughput and faster learning. The policy training is implemented using SKRL reinforcement learning library [32], which provides an integration with Isaac Lab. We model the DSAM system in the training environment. Apart from modeling the control architecture (Section 4.1) with low-level controllers, we also model filters and first-order dynamics of rotors.
PPO offers control over multiple aspects of the algorithm, making hyperparameter selection challenging. We begin with the hyperparameters used in [30] as a starting point.
@@
To achieve a working policy, we train for 2 billion environment steps ($4096\times 500,000$) that take 5.5 hours on a consumer-grade GPU (NVIDIA RTX 2080 Ti).
@@ 5.1 Experimental Setup
The policy trained using the proposed method is deployed on our 2-DoF DSAM, with a total weight of 860g. The moving part (arm plus gripper) weighs 75 g, actuated by two Dynamixel-XC330 motors [29] with embedded PID controllers following joint commands. Our control architecture is developed upon Agilicious [9], a modular software framework designed for agile quadrotor flight. The control stack entirely runs on the DSAM’s onboard CPU (Raspberry Pi 5), including onboard inference using ONNX Runtime [7]. This results in an average inference time of 0.18 ms.
Experiments are conducted in a 13 $\times$ 7 m facility equipped with a motion capture system that provides ground-truth pose measurements at 100 Hz, which is then fused with an EKF running onboard at 300 Hz to obtain the state estimates. A safety controller that only controls the quadrotor base motion [34] is used for take-off and handles workspace violations by immediately bringing the system to hover state. After take-off, control is switched from the safety controller to the learned policy that actuates the quadrotor to reach the commanded end-effector pose.
@@ 5.5 Path following
As seen in Table 3, the RMSE for the figure-8 path is noticeably larger than for the straight line, reflecting the higher demands of multi-axis coordination compared to single-axis motion. The tracking accuracy is further limited because the policy is conditioned only on the immediate target pose and does not explicitly account for higher-order derivatives such as velocity, acceleration, or jerk. Consequently, the policy was not trained specifically for accurate path tracking but rather for pose regulation. Future work could address this by augmenting the observation space with future waypoints or target velocities, which have been shown to improve path-tracking performance in related learning-based control frameworks (e.g.,  [11]). Despite these limitations, the policy is able to follow simple paths while maintaining orientation, demonstrating whole-body coordination emerging from pose control alone.
@@ 6.1.3 Body rate of the quadrotor base
We noticed that including body rate in the observation space is also critical for training for DSAM, which is particularly different from previous works on quadrotors [33], and AMs with a fully-actuated base [6]. We argue that including quadrotor angular rates in the observation significantly improves policy training, as they allow the policy to indirectly infer the effect of torque disturbances induced by the arm. This additional information helps the policy learn the coupled base–arm dynamics especially for DSAM with an underactuated unidirectional base, which is, on the other hand, less critical for omnidirectional platforms or simple quadrotors.
@@ 6.2 Inner-Loop Controller Ablations
While incorporating an inner-loop is a common practice in RL-based control to achieve reliable sim-to-real transfer, determining the appropriate level of control abstraction is non-trivial. In particular, for the quadrotor base of DSAM, we replace the attitude controller with a simple proportional controller that maps the body-rate command to the desired body angular acceleration, which is then passed to the low-level INDI controller. Under this setup, we evaluate the traditional “collective-thrust plus body rate” (CTBR) output of the RL policy for quadrotor flight, following prior work [33]. INDI is retained, as it plays a critical role in handling disturbances, as highlighted in previous studies [34, 8]. The training curves in Fig. 6 show that our setup, where the attitude controller accepts reference acceleration (or equivalently attitude) from the RL policy, achieves consistently higher rewards in both orientation and position tracking.
@@ 7 Conclusion
8 Acknowledgements
This work was supported by the Dutch Research Council (NWO) through the Veni talent Programme (grant no. 20256, accurate aerial Manipulation). We thank Jack Zeng for insightful discussions and for sharing low-level controller templates. We also thank Maurits Pfaff and Kseniia Khomenko for building the 2-DoF DSAM and for their support during experiments.